please help me solve the following problem.
5(t+3)+9=3(t-2)+6

Answer by stanbon(75887)

5(t+3)+9=3(t-2)+6
---
5t+15+9 = 3t-6+6
----
5t+24 = 3t
2t = -24
t = -12
========
Cheers,
Stan H.